# Compact manifolds of dimension $n\geq 12$ with positive isotropic curvature

### v21

@@ 1. $R(x_{k},t_{k})\geq Q_{k}$;
Proof.  See the proof of Theorem 11.2 in [20]. Let $(\mathcal{O},g_{0})$ be as in the assumption. Suppose $|\text{Rm}(g_{0})|\leq K$. We run the Ricci flow starting with $(\mathcal{O},g_{0})$. Let $\hat{T}=\frac{n}{2\inf_{x\in\mathcal{O}}R(x,0)}$. By Theorem 1.2 in [20] and (a version) of Hamilton’s maximum principle, there is a continuous family of closed, convex, $O(n)$-invariant sets $\{\mathcal{F}_{t}~{}|~{}t\in[0,\hat{T}]\}$ in the vector space $\mathcal{C}_{B}(\mathbb{R}^{n})$ of algebraic curvature tensors in dimension $n$ such that the curvature tensor of $(\mathcal{O},g_{0})$ lies in $\mathcal{F}_{0}$, the family is invariant under the Hamilton ODE $\frac{d}{dt}\text{Rm}=Q(\text{Rm})$ and satisfies the inclusion relation (3.1) for some suitable $(f,\theta)$ (as mentioned before), and the curvature tensor of the solution (to this smooth Ricci flow) at time $t$ ($<\hat{T}$) lies in the set $\mathcal{F}_{t}$, hence also satisfies the pinching assumption. Choose $\hat{r}$ and $\hat{\delta}$ as in Proposition 3.9. We may assume that $R_{g_{0}}(x)<(2\hat{r})^{-2}$ for any $x\in\mathcal{O}$ and that for $\delta=\hat{\delta}$ the conclusion of (an orbifold version of) Proposition 8.2 in [20] holds. We may also assume that $\hat{\delta}\leq\tilde{\delta}(\hat{r})$, where $\tilde{\delta}$ is the function in Proposition 3.8. We can choose $\alpha_{1}>0$ such that $(2\hat{r})^{-2}<\max_{(x,t)\in\mathcal{O}\times[0,\alpha_{1}]}R(x,t)<\hat{r}^{-2}$. Then $(\mathcal{O},g(t))$, $t\in[0,\alpha_{1}]$, satisfies (vacuously) the canonical neighborhood assumption $(CN)_{\hat{r}}$ with $(4\varepsilon,4C)$-control. By Proposition 3.9, the flow satisfies $(CN)_{2\hat{r}}$ with $(2\varepsilon,2C)$-control on the time interval $[0,\alpha_{1}]$. By continuity, there exists a positive number $\alpha_{2}$ such that the solution satisfies the canonical neighborhood assumption $(CN)_{\hat{r}}$ with $(4\varepsilon,4C)$-control for $t\in[0,\alpha_{1}+\alpha_{2}]$; cf. for example, Lemmas 5.3.2 and 5.3.3 in [6], Proposition 9.79 in [67], Claim 1 on p. 1229 of [50], and Lemma 5.3. By Proposition 3.9, in fact the solution satisfies the canonical neighborhood assumption $(CN)_{2\hat{r}}$ with $(2\varepsilon,2C)$-control for $t\in[0,\alpha_{1}+\alpha_{2}]$.
@@
To understand the topology of the total space $Y$ of an orbifiber bundle over $\mathbb{S}^{1}$ or $\mathcal{I}$ with generic fiber diffeomorphic to $\mathbb{S}^{n-1}/\Gamma$, we construct its universal covering as follows: First note that if $Y$ is diffeomorphic to (the total space of) an orbifiber bundle over $\mathcal{I}$ with generic fiber diffeomorphic to $\mathbb{S}^{n-1}/\Gamma$, we can take a 2-fold covering to get a $\mathbb{S}^{n-1}/\Gamma$-bundle over $\mathbb{S}^{1}$; cf. for example Section 2.7 of [37]. Actually this 2-fold covering can be constructed by pulling back the orbifiber bundle over $\mathcal{I}$ via the 2-fold orbifold covering $\mathbb{S}^{1}\rightarrow\mathcal{I}$. (By the way, the 2-fold orbifold covering $\mathbb{S}^{1}\rightarrow\mathcal{I}$ is a good map in the sense of Chen-Ruan [32]. Moreover, the pull-back bundle can also be viewed as an orbifold fiber product; cf. for example [32].) For a $\mathbb{S}^{n-1}/\Gamma$-bundle over $\mathbb{S}^{1}$, we can first construct a covering by pulling back this bundle using the universal covering $\mathbb{R}\rightarrow\mathbb{S}^{1}$, thus we get a $\mathbb{S}^{n-1}/\Gamma$-bundle over $\mathbb{R}$, which must be trivial. Then we see that the universal covering of $Y$ is diffeomorphic to $\mathbb{S}^{n-1}\times\mathbb{R}$. So both the manifold $(\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus B_{1})\cup_{h}(\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle\setminus B_{2})$ appearing in the proof of Proposition 3.2 and $Y_{j}$ above are diffeomorphic to quotients of $\mathbb{S}^{n-1}\times\mathbb{R}$ by the corresponding deck transformations, which are certain diffeomorphisms.
@@
5 Appendix: gluing and strengthening of $\varepsilon$-necks
@@
Let $\mathcal{O}$ be a $n$-dimensional orbifold with at most isolated singularities, $U$ an open subset, and $\Gamma$ as above. If $U$ is diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$, we call it a topological neck. As in [29], we divide topological caps into several types. If $U$ is diffeomorphic to $\mathbb{R}^{n}$ or $C_{\Gamma}^{\sigma}$, we call it a smooth cap; if $U$ is diffeomorphic to $\mathbb{R}^{n}//\Gamma$ with $|\Gamma|\geq 2$, we call it an orbifold cap of type I, and denote it by $C_{\Gamma}$; if $U$ is diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus\bar{B}$, we call it an orbifold cap of type II.
@@
For $1\leq k\leq[\varepsilon^{-1}]$, by Lemma 5.1 we have
@@
The following result should also be well-known; cf. for example Proposition 9.79 in [67]. It is used in the proof of Proposition 3.9 and Lemma 5.8.
Lemma 5.3.
@@
As in the proof of Lemma 5.2, we have
@@
For $1\leq k\leq[\varepsilon^{\prime-1}]$, by Lemma 5.1, we have
@@
as in the proof of Lemma 5.2, where $\nabla_{\psi_{1}^{*}g_{2}}$ is the Levi-Civita connection of the metric $\psi_{1}^{*}g_{2}$, and here and below we use $C$ to denote various constants which may be different in different places, and
@@
Let $n\geq 4$. There exists a positive constant $\varepsilon_{1}=\varepsilon_{1}(n)$ with the following property. Let $(\mathcal{O},g)$ be a Riemannian orbifold of dimension $n$ with at most isolated singularities, $0<\varepsilon\leq 2\varepsilon_{1}$, and $N_{i}$ be an $\varepsilon$-neck centered at a point $x_{i}$ in $\mathcal{O}$ given by a diffeomorphism $\psi_{i}:\mathbb{S}^{n-1}/\Gamma_{i}\times(-\varepsilon^{-1},\varepsilon^{-1})\rightarrow N_{i}$, $i=1,2$, such that $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[-0.9\varepsilon^{-1},0.9\varepsilon^{-1}])\cap N_{2}\neq\emptyset$. Then $\Gamma_{1}$ and $\Gamma_{2}$ are conjugate in $O(n)$, and a cross section of the neck $N_{1}$ (that is the image $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{i}\times\{z\})$ for some $z\in(-\varepsilon^{-1},\varepsilon^{-1})$) and a cross section of the neck $N_{2}$ cobound a compact domain in $N_{1}\cup N_{2}$ diffeomorphic to the product $\mathbb{S}^{n-1}/\Gamma_{1}\times[0,1]$ if they are disjoint.
@@
Proof.   We use a Morse theoretical argument. Note that Lemma A.4 and item (3) of Proposition A.11 in [67] (there is also a statement in the proof of Lemma G1.4 in [48] which is similar to Lemma A.4 in [67]) can be extended to the case of $n$-dimensional $\varepsilon$-necks; see also the proof of Proposition 1.16 and Corollary 3.7 in [40]. (For example, to extend Lemma A.4 in [67] to our situation we can combine the argument in the proof of Lemma A.4 in [67] with techniques from the covering space theory.) Since $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[-0.9\varepsilon^{-1},0.9\varepsilon^{-1}])\cap N_{2}\neq\emptyset$, if $\varepsilon_{1}=\varepsilon_{1}(n)$ is sufficient small and $0<\varepsilon\leq 2\varepsilon_{1}$, without loss of generality we may assume that $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[0.901\varepsilon^{-1},0.9999\varepsilon^{-1}])\subset N_{2}$, and that there exists a cross section, say $S$, of the neck $N_{2}$ such that $S\subset\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[0.994\varepsilon^{-1},0.995\varepsilon^{-1}])$. We rescale the metric $g$ by a constant factor such that after rescaling the scalar curvature on $N_{1}\cap N_{2}$ is close to 1. Let $p_{2}:\mathbb{S}^{n-1}/\Gamma_{1}\times\mathbb{R}\rightarrow\mathbb{R}$ and $p_{2}^{\prime}:\mathbb{S}^{n-1}/\Gamma_{2}\times\mathbb{R}\rightarrow\mathbb{R}$ be the natural projections onto the second factor. The following estimates are for quantities computed w.r.t. the rescaled metric (on $N_{1}\cap N_{2}$) under the assumption that $0<\varepsilon\leq 2\varepsilon_{1}(n)$ with $\varepsilon_{1}(n)$ sufficiently small. We have $||\nabla(p_{2}\circ\psi_{1}^{-1})|-1|\ll 1$ and $||\nabla(p_{2}^{\prime}\circ\psi_{2}^{-1})|-1|\ll 1$ on $N_{1}\cap N_{2}$. Using the $n$-dimensional extension of item (3) of Proposition A.11 in [67] we see that after replacing $\psi_{2}$ by $\psi_{2}$ precomposed with a reflection (that is reversing the $\varepsilon$-neck structure of $N_{2}$) if necessary, the restriction of the gradient vector field $\nabla(p_{2}\circ\psi_{1}^{-1})$ to $N_{1}\cap N_{2}$ and the restriction of $\nabla(p_{2}^{\prime}\circ\psi_{2}^{-1})$ to $N_{1}\cap N_{2}$ are almost parallel to each other with the same direction; one can also see this by using the fact that both $\nabla(p_{2}\circ\psi_{1}^{-1})$ and $\nabla(p_{2}^{\prime}\circ\psi_{2}^{-1})$ are very close to the eigenspace of Ric (viewed as a (1,1)-tensor) corresponding to the smallest eigenvalue at each point of $N_{1}\cap N_{2}$ (the fact in turn follows from the closeness of $g|_{N_{1}\cap N_{2}}$ to the metrics on the corresponding parts of the two quotients $\mathbb{S}^{n-1}/\Gamma_{i}\times(-\varepsilon^{-1},\varepsilon^{-1})$ ($i=1,2$) of the round cylinder). Cf. the proof of Lemma 5.4.5 in [2]. So $|\nabla(p_{2}\circ\psi_{1}^{-1}-p_{2}^{\prime}\circ\psi_{2}^{-1})|\ll 1$ on $N_{1}\cap N_{2}$. After a translation of the parameter for the parametrization $\psi_{2}$ (the shifted parametrization will still be denoted by $\psi_{2}$) we may assume that $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times\{0.95\varepsilon^{-1}\})\cap\psi_{2}(\mathbb{S}^{n-1}/\Gamma_{2}\times\{0.95\varepsilon^{-1}\})\neq\emptyset$.
@@
Let $W=\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[0.901\varepsilon^{-1},0.9999\varepsilon^{-1}])$, $\rho:W\rightarrow[0,1]$ be a smooth cutoff function which is 1 on $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[0.905\varepsilon^{-1},0.996\varepsilon^{-1}])$ and 0 outside $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times(0.902\varepsilon^{-1},0.999\varepsilon^{-1}))$, and $\zeta:W\rightarrow[0,1]$ be a suitable smooth function which is 1 on $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[0.901\varepsilon^{-1},0.908\varepsilon^{-1}])$ and 0 on $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[0.993\varepsilon^{-1},0.9999\varepsilon^{-1}])$ with $|\nabla\zeta|<20\varepsilon$. We construct a smooth function $h=\zeta p_{2}\circ\psi_{1}^{-1}|_{W}+(1-\zeta)p_{2}^{\prime}\circ\psi_{2}^{-1}|_{W}$ on $W$. Note that $|p_{2}\circ\psi_{1}^{-1}-p_{2}^{\prime}\circ\psi_{2}^{-1}|\cdot|\nabla\zeta|\ll 1$ on $W$. So if $0<\varepsilon\leq 2\varepsilon_{1}(n)$ with $\varepsilon_{1}(n)$ sufficiently small, the gradient vector field $\nabla h\neq 0$ everywhere on $W$. Consider the vector field $X=\frac{\rho}{|\nabla h|^{2}}\nabla h$ on $W$. (Cf. Milnor [65].) Using the flow generated by the vector field $X$ we see that the hypersurfaces $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times\{0.907\varepsilon^{-1}\})$ and $S$ cobound a compact domain in $W$ which is diffeomorphic to the product $\mathbb{S}^{n-1}/\Gamma_{1}\times[0,1]$. (Compare the proof of item (5) of Proposition A.11 in [67], the proof of Theorem G1.1 in [48], the last paragraph in the proof of Theorem 3.9 in [29], and the second to last paragraph in the proof of Proposition 3.6 in [51].) In particular, the two hypersurfaces are diffeomorphic. By [39] and Lemma 2.5.6 in [82], $\Gamma_{1}$ and $\Gamma_{2}$ are conjugate in $O(n)$. Now by using Lemma 3.1 we see that a cross section of the neck $N_{1}$ and a cross section of the neck $N_{2}$ cobound a compact domain in $N_{1}\cup N_{2}$ diffeomorphic to the product $\mathbb{S}^{n-1}/\Gamma_{1}\times[0,1]$ if they are disjoint. $\Box$
@@
Remark. In the special case that both $\Gamma_{1}$ and $\Gamma_{2}$ are trivial and the dimension $n\geq 5$, we have a more topological argument to show the conclusion of Proposition 5.4 along the lines of proof of Lemma 3.2.2 in [6]; see the proof of Claim 1 in the proof of Proposition 2.5.
@@
The following result is an application of Proposition 5.4 (and the Cerf-Palais disk theorem). It is an extension of Lemma 3.2.8 in [6], and is used in the proof of Proposition 3.3.
Lemma 5.5.
@@
(cf. Lemma 3.2.8 in [6]) Let $n\geq 4$, and $0<\varepsilon\leq 2\varepsilon_{1}$, where $\varepsilon_{1}$ is as in Proposition 5.4. Suppose that $M$ is a connected Riemannian manifold of dimension $n$, and $\gamma:[0,1)\rightarrow M$ is a minimal geodesic such that $\gamma(s)$ is a center of an $\varepsilon$-neck in $M$ diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$ for each $s\in[0,1)$, and $R(\gamma(s))\rightarrow\infty$ as $s\rightarrow 1$. Then there exists an $\varepsilon$-horn in $M$ diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$ which contains the open end of the image of $\gamma$.

### v22

@@ 1. $R(x_{k},t_{k})\geq Q_{k}$;
Proof.  See the proof of Theorem 11.2 in [21]. Let $(\mathcal{O},g_{0})$ be as in the assumption. Suppose $|\text{Rm}(g_{0})|\leq K$. We run the Ricci flow starting with $(\mathcal{O},g_{0})$. Let $\hat{T}=\frac{n}{2\inf_{x\in\mathcal{O}}R(x,0)}$. By Theorem 1.2 in [21] and (a version) of Hamilton’s maximum principle, there is a continuous family of closed, convex, $O(n)$-invariant sets $\{\mathcal{F}_{t}~{}|~{}t\in[0,\hat{T}]\}$ in the vector space $\mathcal{C}_{B}(\mathbb{R}^{n})$ of algebraic curvature tensors in dimension $n$ such that the curvature tensor of $(\mathcal{O},g_{0})$ lies in $\mathcal{F}_{0}$, the family is invariant under the Hamilton ODE $\frac{d}{dt}\text{Rm}=Q(\text{Rm})$ and satisfies the inclusion relation (3.1) for some suitable $(f,\theta)$ (as mentioned before), and the curvature tensor of the solution (to this smooth Ricci flow) at time $t$ ($<\hat{T}$) lies in the set $\mathcal{F}_{t}$, hence also satisfies the pinching assumption. Choose $\hat{r}$ and $\hat{\delta}$ as in Proposition 3.9. We may assume that $R_{g_{0}}(x)<(2\hat{r})^{-2}$ for any $x\in\mathcal{O}$ and that for $\delta=\hat{\delta}$ the conclusion of (an orbifold version of) Proposition 8.2 in [21] holds. We may also assume that $\hat{\delta}\leq\tilde{\delta}(\hat{r})$, where $\tilde{\delta}$ is the function in Proposition 3.8. We can choose $\alpha_{1}>0$ such that $(2\hat{r})^{-2}<\max_{(x,t)\in\mathcal{O}\times[0,\alpha_{1}]}R(x,t)<\hat{r}^{-2}$. Then $(\mathcal{O},g(t))$, $t\in[0,\alpha_{1}]$, satisfies (vacuously) the canonical neighborhood assumption $(CN)_{\hat{r}}$ with $(4\varepsilon,4C)$-control. By Proposition 3.9, the flow satisfies $(CN)_{2\hat{r}}$ with $(2\varepsilon,2C)$-control on the time interval $[0,\alpha_{1}]$. By continuity, there exists a positive number $\alpha_{2}$ such that the solution satisfies the canonical neighborhood assumption $(CN)_{\hat{r}}$ with $(4\varepsilon,4C)$-control for $t\in[0,\alpha_{1}+\alpha_{2}]$; cf. for example, Lemmas 5.3.2 and 5.3.3 in [6], Proposition 9.79 in [70], Claim 1 on p. 1229 of [53], and Lemma A.3. By Proposition 3.9, in fact the solution satisfies the canonical neighborhood assumption $(CN)_{2\hat{r}}$ with $(2\varepsilon,2C)$-control for $t\in[0,\alpha_{1}+\alpha_{2}]$.
@@
To understand the topology of the total space $Y$ of an orbifiber bundle over $\mathbb{S}^{1}$ or $\mathcal{I}$ with generic fiber diffeomorphic to $\mathbb{S}^{n-1}/\Gamma$, we construct its universal covering as follows: First note that if $Y$ is diffeomorphic to (the total space of) an orbifiber bundle over $\mathcal{I}$ with generic fiber diffeomorphic to $\mathbb{S}^{n-1}/\Gamma$, we can take a 2-fold covering to get a $\mathbb{S}^{n-1}/\Gamma$-bundle over $\mathbb{S}^{1}$; cf. for example Section 2.7 of [40]. Actually this 2-fold covering can be constructed by pulling back the orbifiber bundle over $\mathcal{I}$ via the 2-fold orbifold covering $\mathbb{S}^{1}\rightarrow\mathcal{I}$. (By the way, although in a local model (uniformizing system) of an orbifold considered in [34] any element of the group is required to either act trivially or have fixed-point set of codimension at least two (see p. 65 of [34]), one can extend their notions of orbifolds and good maps slightly so that the 2-fold orbifold covering $\mathbb{S}^{1}\rightarrow\mathcal{I}$ is a good map in the extended sense. Moreover, the pull-back bundle can also be viewed as an orbifold fiber product; cf. for example p. 83 in [34] and Chapter 13 in [84].) For a $\mathbb{S}^{n-1}/\Gamma$-bundle over $\mathbb{S}^{1}$, we can first construct a covering by pulling back this bundle using the universal covering $\mathbb{R}\rightarrow\mathbb{S}^{1}$, thus we get a $\mathbb{S}^{n-1}/\Gamma$-bundle over $\mathbb{R}$, which must be trivial. Then we see that the universal covering of $Y$ is diffeomorphic to $\mathbb{S}^{n-1}\times\mathbb{R}$. So both the manifold $(\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus B_{1})\cup_{h}(\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle\setminus B_{2})$ appearing in the proof of Proposition 3.2 and $Y_{j}$ above are diffeomorphic to quotients of $\mathbb{S}^{n-1}\times\mathbb{R}$ by the corresponding deck transformations, which are certain diffeomorphisms.
@@
Appendix A Appendix: gluing and strengthening of $\varepsilon$-necks
@@
Let $\mathcal{O}$ be an $n$-dimensional orbifold with at most isolated singularities, $U$ an open subset, and $\Gamma$ as above. If $U$ is diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$, we call it a topological neck. As in [31], we divide topological caps into several types. If $U$ is diffeomorphic to $\mathbb{R}^{n}$ or $C_{\Gamma}^{\sigma}$, we call it a smooth cap; if $U$ is diffeomorphic to $\mathbb{R}^{n}//\Gamma$ with $|\Gamma|\geq 2$, we call it an orbifold cap of type I, and denote it by $C_{\Gamma}$; if $U$ is diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus\bar{B}$, we call it an orbifold cap of type II.
@@
For $1\leq k\leq[\varepsilon^{-1}]$, by Lemma A.1 we have
@@
The following result should also be well-known; cf. for example Proposition 9.79 in [70]. It is used in the proof of Proposition 3.9 and Lemma A.8.
Lemma A.3.
@@
As in the proof of Lemma A.2, we have
@@
For $1\leq k\leq[\varepsilon^{\prime-1}]$, by Lemma A.1, we have
@@
as in the proof of Lemma A.2, where $\nabla_{\psi_{1}^{*}g_{2}}$ is the Levi-Civita connection of the metric $\psi_{1}^{*}g_{2}$, and here and below we use $C$ to denote various constants which may be different in different places, and
@@
Let $n\geq 4$. There exists a positive constant $\varepsilon_{1}=\varepsilon_{1}(n)$ with the following property. Let $(\mathcal{O},g)$ be a Riemannian orbifold of dimension $n$ with at most isolated singularities, $0<\varepsilon\leq 2\varepsilon_{1}$, and $N_{i}$ be an $\varepsilon$-neck centered at a point $x_{i}$ in $\mathcal{O}$ given by a diffeomorphism $\psi_{i}:\mathbb{S}^{n-1}/\Gamma_{i}\times(-\varepsilon^{-1},\varepsilon^{-1})\rightarrow N_{i}$, $i=1,2$, such that $N_{1}\cap\psi_{2}(\mathbb{S}^{n-1}/\Gamma_{2}\times[-0.9\varepsilon^{-1},0.9\varepsilon^{-1}])\neq\emptyset$. Then $\Gamma_{1}$ and $\Gamma_{2}$ are conjugate in $O(n)$, and there is a cross-section of the neck $N_{1}$ (that is the image $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times\{z\})$ for some $z\in(-\varepsilon^{-1},\varepsilon^{-1})$) and a cross-section of the neck $N_{2}$ which cobound a compact domain in $N_{1}\cup N_{2}$ diffeomorphic to the product $\mathbb{S}^{n-1}/\Gamma_{1}\times[0,1]$.
@@
Proof.   We use a Morse theoretical argument. Note that Lemma A.4 in [70] (there is also a statement in the proof of Lemma G1.4 in [51] which is similar to Lemma A.4 in [70]) can be extended to the case of $n$-dimensional $\varepsilon$-necks. (For example, to extend Lemma A.4 in [70] to our situation we can combine the argument in the proof of Lemma A.4 in [70] with techniques from the covering space theory.) Since $N_{1}\cap\psi_{2}(\mathbb{S}^{n-1}/\Gamma_{2}\times[-0.9\varepsilon^{-1},0.9\varepsilon^{-1}])\neq\emptyset$, if $\varepsilon_{1}=\varepsilon_{1}(n)$ is sufficiently small and $0<\varepsilon\leq 2\varepsilon_{1}$, the ratio of the scales of the two $\varepsilon$-necks $N_{1}$ and $N_{2}$ is very close to 1, and there exist numbers $a,b\in(-\varepsilon^{-1},\varepsilon^{-1})$ with $a<b$ and $b-a\geq 0.099\varepsilon^{-1}$ such that $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[a,b])\subset N_{2}$. Then one sees that there exists a cross section, say $S$, of the neck $N_{2}$ such that $S\subset\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[a+0.093\varepsilon^{-1},a+0.094\varepsilon^{-1}])$. We rescale the metric $g$ by a constant factor such that after rescaling the scalar curvature on $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[a,b])$ is close to 1. Let $p_{2}:\mathbb{S}^{n-1}/\Gamma_{1}\times\mathbb{R}\rightarrow\mathbb{R}$ and $p_{2}^{\prime}:\mathbb{S}^{n-1}/\Gamma_{2}\times\mathbb{R}\rightarrow\mathbb{R}$ be the natural projections onto the second factor. The following estimates are for quantities computed w.r.t. the rescaled metric (on $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[a,b])$) under the assumption that $0<\varepsilon\leq 2\varepsilon_{1}(n)$ with $\varepsilon_{1}(n)$ sufficiently small. We have $||\nabla(p_{2}\circ\psi_{1}^{-1})|-1|\ll 1$ and $||\nabla(p_{2}^{\prime}\circ\psi_{2}^{-1})|-1|\ll 1$ on $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[a,b])$. Using the $n$-dimensional extension of Lemma A.4 in [70] we see that after replacing $\psi_{2}$ by $\psi_{2}$ precomposed with a reflection (that is reversing the direction of the $\varepsilon$-neck structure of $N_{2}$) if necessary, the directions of the two vector fields, that is, the gradient vector fields $\nabla(p_{2}\circ\psi_{1}^{-1})$ and $\nabla(p_{2}^{\prime}\circ\psi_{2}^{-1})$ both restricted to the region $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[a,b])$, are very close to each other; one can also see this by using the fact that both $\nabla(p_{2}\circ\psi_{1}^{-1})$ and $\nabla(p_{2}^{\prime}\circ\psi_{2}^{-1})$ (restricted to the region $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[a,b])$), are very close to the eigenspace of Ric (viewed as a (1,1)-tensor) corresponding to the smallest eigenvalue at each point of $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[a,b])$ (this fact in turn follows from the closeness of the metric $g$ (restricted to $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[a,b])$) to the metrics on the corresponding parts of the two quotients $\mathbb{S}^{n-1}/\Gamma_{i}\times(-\varepsilon^{-1},\varepsilon^{-1})$ ($i=1,2$) of the round cylinder). Cf. the proof of Lemma 5.4.5 in [2]. So $|\nabla(p_{2}\circ\psi_{1}^{-1}-p_{2}^{\prime}\circ\psi_{2}^{-1})|\ll 1$ on $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[a,b])$. After translations of the parameters for the parametrizations $\psi_{i}$ (the shifted parametrizations will still be denoted by $\psi_{i}$) we may assume that $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[0.9\varepsilon^{-1},0.999\varepsilon^{-1}])\subset N_{2}$ and $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times\{0.95\varepsilon^{-1}\})\cap\psi_{2}(\mathbb{S}^{n-1}/\Gamma_{2}\times\{0.95\varepsilon^{-1}\})\neq\emptyset$.
@@
Let $W=\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[0.9\varepsilon^{-1},0.999\varepsilon^{-1}])$, $\rho:W\rightarrow[0,1]$ be a smooth cutoff function which is 1 on $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[0.904\varepsilon^{-1},0.995\varepsilon^{-1}])$ and 0 outside $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times(0.901\varepsilon^{-1},0.998\varepsilon^{-1}))$, and $\zeta:W\rightarrow[0,1]$ be a suitable smooth function which is 1 on $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[0.9\varepsilon^{-1},0.907\varepsilon^{-1}])$ and 0 on $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times[0.992\varepsilon^{-1},0.999\varepsilon^{-1}])$ with $|\nabla\zeta|<20\varepsilon$. We construct a smooth function $h=\zeta p_{2}\circ\psi_{1}^{-1}|_{W}+(1-\zeta)p_{2}^{\prime}\circ\psi_{2}^{-1}|_{W}$ on $W$. Note that $|p_{2}\circ\psi_{1}^{-1}-p_{2}^{\prime}\circ\psi_{2}^{-1}|\cdot|\nabla\zeta|\ll 1$ on $W$. So if $0<\varepsilon\leq 2\varepsilon_{1}(n)$ with $\varepsilon_{1}(n)$ sufficiently small, the gradient vector field $\nabla h\neq 0$ everywhere on $W$. Consider the vector field $X=\frac{\rho}{|\nabla h|^{2}}\nabla h$ on $W$. (Cf. Milnor [68].) Using the flow generated by the vector field $X$ we see that the hypersurfaces $\psi_{1}(\mathbb{S}^{n-1}/\Gamma_{1}\times\{0.906\varepsilon^{-1}\})$ and $S$ (defined in the last paragraph) cobound a compact domain in $W$ which is diffeomorphic to the product $\mathbb{S}^{n-1}/\Gamma_{1}\times[0,1]$. (Compare the proof of item (5) of Proposition A.11 in [70], the proof of Theorem G1.1 in [51], the last paragraph in the proof of Theorem 3.9 in [31], and the second to last paragraph in the proof of Proposition 3.6 in [54].) In particular, the two hypersurfaces are diffeomorphic. By [42] and Lemma 2.5.6 in [88], $\Gamma_{1}$ and $\Gamma_{2}$ are conjugate in $O(n)$. $\Box$
@@
Remark. In the special case that both $\Gamma_{1}$ and $\Gamma_{2}$ are trivial and the dimension $n\geq 5$, we have a more topological argument to show the conclusion of Proposition A.4 along the lines of proof of Lemma 3.2.2 in [6]; see the proof of Claim 1 in the proof of Proposition 2.5.
@@
The following result is an application of Proposition A.4 (and the Cerf-Palais disk theorem). It is an extension of Lemma 3.2.8 in [6], and is used in the proof of Proposition 3.3.
Lemma A.5.
@@
(cf. Lemma 3.2.8 in [6]) Let $n\geq 4$, and $0<\varepsilon\leq 2\varepsilon_{1}$, where $\varepsilon_{1}$ is as in Proposition A.4. Suppose that $M$ is a connected Riemannian manifold of dimension $n$, and $\gamma:[0,1)\rightarrow M$ is a minimal geodesic such that $\gamma(s)$ is a center of an $\varepsilon$-neck in $M$ diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$ for each $s\in[0,1)$, and $R(\gamma(s))\rightarrow\infty$ as $s\rightarrow 1$. Then there exists an $\varepsilon$-horn in $M$ diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$ which contains the open end of the image of $\gamma$.